Please help me with this equation: A restaurant manager buys 100lb of hambure and 50lb of steak for a total cost of $490. A second purchase at the same prices, includes 150lb of hambureger and 100lb of steak. Total cost is 860. Find the price of 1lb of steak.
Answer by jorel1380(3719)

100m+50n=490
150m+100n=860
200m+100n=980
50m=980-860=120
m=120/50=$2.40 per lb of hamburger
n=$5.00 per lb. of steak..